solve(3x)^2-12√(3x)+27=0 using the quadratic equation.
explain the steps please?

Answer by stanbon(75887)

solve(3x)^2-12√(3x)+27=0 using the quadratic equation.
-----
It is not a quadratic the way you have posted it.
----
The following would be the quadratic form:
(3x)^2 - 12(3x) + 27 = 0
----
Letting w = 3x you get:
w^2 - 12w + 27 = 0
(w-9)(w-3) = 0
w = 9 or w = 3
-----
Solving for "x":
3x=9 or 3x=3
x = 3 or x = 1
====================
Cheers,
Stan H.
==============